# INACIA: Integrating Large Language Models in Brazilian Audit Courts: Opportunities and Challenges

### 10 Jan 2024 v1

@@ Abstract.
This paper introduces INACIA (Instrução Assistida com Inteligência Artificial), a groundbreaking system designed to integrate Large Language Models (LLMs) into the operational framework of Brazilian Federal Court of Accounts (TCU). The system automates various stages of case analysis, including basic information extraction, admissibility examination, Periculum in mora and Fumus boni iuris analyses, and recommendations generation. Through a series of experiments, we demonstrate INACIA’s potential in extracting relevant information from case documents, evaluating its legal plausibility, and generating judicial recommendations. Utilizing a validation dataset alongside LLMs, our evaluation methodology presents an innovative approach to assessing system performance, correlating highly with human judgment. The results highlight INACIA’s proficiency in handling complex legal tasks, indicating its suitability for augmenting efficiency and judicial fairness within legal systems. The paper also discusses potential enhancements and future applications, positioning INACIA as a model for worldwide AI integration in legal domains.
@@ 1. Introduction
In this paper, we introduce a groundbreaking application of LLMs for administrative case processing and autonomous decision-making developed jointly by the Brazilian Federal Court of Accounts (TCU, or Tribunal de Contas da União, in Portuguese) and experts at several Brazilian research organizations. INACIA (Instrução Assistida com Inteligência Artificial, in Portuguese or Instruction Assisted with Artificial Intelligence) is a system that analyzes and issues decisions [^1] These decisions are somewhat similar to expert opinions but rather than being offered by external experts, these are part of the responsibilities of staff auditors at TCU. They are “decisions” in the sense that they can be adopted in full and are binding once staff auditors and division heads sign off on the recommended course of action. for a subset of cases under TCU’s responsibility. Its inputs are case filings; outputs are (official) court decision documents and tabular data required for the court’s record-keeping activities. INACIA uses a combination of LLMs and search algorithms to extract information from case documents and to reason through the extracted information.
The main challenge in developing systems like INACIA is evaluating their performance. How can we ensure the quality of the produced decisions and extracted data? To address this challenge, we construct a validation dataset from TCU’s case filings and propose a methodology that uses LLMs to evaluate the system’s outputs. To our knowledge, this is the first evaluation of this kind in the legal domain. We demonstrate that our methodology is highly correlated with human evaluation and can be used to evaluate the system’s performance. Our results demonstrate that INACIA can produce high-quality decisions but also forewarn the need for further research to improve the system’s performance.
@@ 2.1. Role and Responsibilities
In most contemporary democracies, a network of institutions exists to regulate, investigate, and potentially sanction individuals occupying elected and non-elected office for improperly using public funds (Speck, 2011). Despite corruption often dominating media attention, mismanagement of public resources makes up the majority of audit findings across Brazil (Ferraz and Finan, 2008). TCU occupies a premier position within the control network of the Brazilian government and is seen as a model institution for auditing practices across the country. [^2] Each state in Brazil also has its state audit court heavily inspired by TCU and closely tied to the Federal Court.
@@ 2.2. Artificial Intelligence and TCU
Artificial Intelligence (AI) technologies are not new to TCU. Over the past five years, court developers have been working on systems to summarize case documents, to efficiently search and retrieve case law information to support decisions, and, more recently, to create an internal ChatGPT-like conversational bot to help staff find all kinds of unstructured, internal information. Earlier in 2023, TCU created its own Artificial Intelligence Center (or Núcleo de Inteligência Artificial, in Portuguese), whose goal is to guide and implement responsible, ethical use of AI in the court.
Recognizing the rapid evolution of AI systems, large language models in particular, TCU issued a call for proposals in early 2022, inviting external vendors to submit proposals to develop an end-to-end system to streamline audit case processing. TCU aims to focus auditors’ expertise on the most challenging tasks of case processing while handing over the more routine tasks to the AI system. This strategy is well-aligned with the evidence on how best to incorporate technology for productivity gains and labor output augmentation (Acemoglu and Autor, 2011).
@@ 2.3. Audit Case Processing
While case proceedings may vary based on the initiation method, case processing is typically a time-intensive task primarily handled by TCU staff. Auditors have to review case filings, confirm whether TCU should hear the case, search and append external evidence of wrongdoing (if necessary), check claims against case law and federal statutes, apply the Court’s jurisprudence and prepare a final document summarizing everything, including the recommended action on the case, to be sent over to a panel of judges for final adjudication. TCU estimates that reviewing a single case requires approximately 30 hours, incurring a cost of approximately \$1,750 per case before a case is even presented to the court’s judicial panel. Given an annual average processing rate of about 2,000 cases and the textual nature of the task, audit case processing is an ideal candidate for an LLM application that would help expedite review while reducing costs for the organization. [^3] For comparison purposes, INACIA processes a case in under 20 minutes and at an initial cost of \$10, subject to further human review. A mere 1 percent reduction in the time required for a human to process a case would suffice to offset the costs associated with INACIA.
2.4. Vision for AI at the Court
TCU is an early adopter of AI technology, and INACIA was one of its major AI endeavors. INACIA will thus shape future AI adoption at the court. The current expectation is that AI will change how TCU handles cases from a sequential, step-by-step approach to a more holistic approach in which filings are evaluated simultaneously and connected to external resources necessary for decision-making from the outset. Effectively, this system would eliminate redundancies and streamline case processing without requiring regulation change.
@@ 3.1. Overview
Figure 1 summarizes INACIA’s main case processing pipeline, showcasing the holistic nature of case analysis. While there is a direct, sequential flow (i.e., solid arrows) of tasks (i.e., blue boxes) needed to produce case decision documents, the dashed lines show the association (or dependency) between tasks. The lighter grey boxes indicate data objects needed to start processing or to be produced as intermediate or final products for other TCU systems. In the next subsections, we describe each element of this pipeline in detail.
@@ 3.2. Inputs
Whenever Tika fails in text extraction (e.g., it reported zero searchable characters after processing the text), INACIA then deploys the second text extraction solution, which relies on Microsoft’s Azure Form Recognizer service for Optical Character Recognition (OCR) (Microsoft, [n. d.]). Form Recognizer will ingest any unsearchable file and return the extracted text in plain format. We also compared Form Recognizer to other industry-leading OCR frameworks (Tesseract (Smith, 2007), and Google Cloud Vision (Google, [n. d.])) using the high-quality and low-cost criteria. While they are similarly priced, Form Recognizer had higher-quality results and richer extraction metadata (e.g., information on the position and content of tables).444 To assess text extraction from either document type, we evaluated results qualitatively. We manually compared extracted elements from case documents, such as the position of tables, columns, and text, allowing for a more nuanced understanding of the strengths and weaknesses of each tool in practical scenarios. We evaluated a sample of documents from TCU’s database containing all cases and their documents between 2019 and 2022. We pulled out a minimum of five documents per year. Within these documents, we had at least one document of easy, medium, and hard extraction, determined by the number of pages and their content as follows: (1) Easy: up to 10-page documents with minimal tables; (2) Medium: up to 25-page documents with structured data or images; (3) Hard: over 25-page documents with structured and unstructured data elements, images, or handwritten content.
@@ 3.3.1. Basic Information Extraction
Figure 2 illustrates the basic information extraction process. It is a structured and systematic process relying on prompt engineering and retrieval-augmented generation (RAG) techniques (Lewis et al., 2020): we will first feed the LLM a structured list of the information we need to extract from case documents (“Pre-defined form” green box). This list was constructed using older cases and their representação containing the information TCU needs to retain for record-keeping purposes 555 “Basic information” examples are the case ID, case type, plaintiff name, plaintiff ID, etc. . We also send over detailed explanations of what each of the variables is. Finally, we append to our combined prompt from above the first two pages of a case’s representação (“Main document” gray box), which should contain most of the information needed.
@@ 3.3.2. Extraction of Allegations and Requests
An important step in our pipeline is to assess whether a case’s claims and supporting evidence are strong. Often, TCU staff can turn down the review of a case when there is weak evidence of wrongdoing when the plaintiff cannot straightforwardly articulate their claims that the defendant has broken the law in some shape or form, or even when the plaintiff does not clearly states their request for TCU action. As with other tasks, extracting allegations and actions requested by plaintiffs is a time-consuming task suited for LLM processing.
@@
The data extracted in this step (allegations and requests) are used in three ways: (1) for the admissibility examination, where INACIA reviews cases for elements that would make such case “of public interest” and whether the claims are strong enough to hear the case; (2) for drafting the final decision document (called “initial instruction” or instrução inicial in Portuguese and cf. Section 3.7); (3) for the Fumus Boni Iuris analysis, which means evaluating a case for the likelihood of success on its merits (cf. Section 3.6).
@@ 3.4. Admissibility Examination
In this task, the agent can compose queries to search for information in case documents and external databases. For example, the agent may want to compose the query quem tem legitimidade para apresentar uma representação ao TCU? (i.e., “who has legitimacy to file a claim to the TCU?”) to search for information about who can file a claim in external databases. If that’s the case, the agent should not be limited to pre-defined queries and keyword-based queries. Therefore, the search engine, such as our implemented solution combining the BM25 algorithm and the monoT5666The monoT5 re-ranker is a neural network that re-ranks the documents based on the query and the document’s content. The re-ranker is a T5 model (Raffel et al., 2020) fine-tuned to indicate whether a document is relevant to a given query. The model produces a score between 0 and 1, indicating the document’s relevance to the query. This score is used to re-rank the documents. re-ranker (Nogueira et al., 2020). The BM25 algorithm is used to retrieve the documents that contain the words in the query. The monoT5 re-ranker is then used to re-rank the top 1,000 documents retrieved by the BM25 algorithm.
@@ 3.5. Periculum in mora Analysis
The analysis, however, is relatively straightforward. We carry out a simple search for exact matches of minuta de contrato (draft contract) in all case documents after a case has been assigned to a TCU staff, except when the case contains fewer than five documents – in which case all documents are considered. This search is essential to flag documents mentioning the draft contract so they are not included in subsequent steps. Two key rules are used to standardize the decision on whether there is Periculum in mora: (1) the existence of an active or signed contract – there is no danger in delay since TCU cannot prevent a contract from coming into force; (2) and (2) the existence of events that prove delay, cancellation, or impediment in contracting – there is no danger in delay because the execution of a contract has been halted. The verification of these rules occurs step-wise. We first verify whether there is any indication that a goods or service provision contract has been signed or is active. The verification is done directly from a keyword search and indirectly by asking the LLM to reason through the documents and verify the existence and status of procurement contracts. The second step is verifying any event delaying, canceling, or suspending a contract from entering into force. Here, we do not perform a direct keyword search for such information; we send documents and prompts straight to the LLM. The outcome of these searches provides an answer regarding the existence of an active contract and any delays or impediments. Finally, the findings are synthesized into a dictionary containing information about the presence of an active contract or delays, along with the document IDs from each search, and a decision on whether there is Periculum in mora, which can be either “rejected” or “accepted”.
It is important to note that the role of the LLM is limited to receiving search results and determining the existence of an active contract or impediments and delays. The LLM does not engage in function calling or actions like the ReAct framework (Yao et al., 2023), where the model defines actions to be taken. The ReAct framework is a paradigm that leverages reasoning and acting within language models to tackle a variety of language, reasoning, and decision-making tasks. The decision-making flows are hard-coded and not subject to the LLM’s discretion. After the decision is made, the result is passed onto the LLM along with a prompt laying out decision rules, enabling the LLM to draft the text that will be included in the final decision document (the “instruction”). This process ensures that the conclusions drawn from the Periculum in mora analysis are accurately reflected in the official documentation, preserving the integrity of TCU’s case review process.
@@ 3.7. Recommendations
The INACIA pipeline concludes with recommendations to the adjudication panel drafted by TCU staff assigned to review each. We call this the “Recommendations” stage. This pipeline task is relatively straightforward and recasts the other outputs from each task to construct a cohesive and coherent decision document to be sent over to the panel of judges who will issue the final ruling on a case. To write case recommendations, INACIA uses the following components:
  • The output from the admissibility examination, i.e., whether the case is admissible and the explanation of such decision.
@@
Each of these inputs is then processed by the LLM, which had been previously instructed with guidelines and examples of recommendations authored by experienced auditors. The LLM generates paragraphs for each recommendations section, ensuring the final text is coherent and accurately reflects the decisions and analyses. The end product is a structured document that is broken down into several sections, each on different aspects of the case, such as the case information, decisions on admissibility, merits, etc.
This modular construction of the final decision document makes sense both from the legal and the computational sense. From the legal perspective, it allows human reviewers to parse out the document into smaller chunks and use/correct them as they see fit. From the computational side, it means we do not have to reprocess everything in later stages if we need a small correction in one of the earlier stages of INACIA. This approach provides efficiency, especially in cases where, despite a potential flaw in one section, the rest of the information remains valid and useful.

### 19 Jan 2024 v2

@@ Abstract.
This paper introduces INACIA (Instrução Assistida com Inteligência Artificial), a groundbreaking system designed to integrate Large Language Models (LLMs) into the operational framework of Brazilian Federal Court of Accounts (TCU). The system automates various stages of case analysis, including basic information extraction, admissibility examination, Periculum in mora and Fumus boni iuris analyses, and recommendations generation. Through a series of experiments, we demonstrate INACIA’s potential in extracting relevant information from case documents, evaluating its legal plausibility, and formulating propositions for judicial decision-making. Utilizing a validation dataset alongside LLMs, our evaluation methodology presents an innovative approach to assessing system performance, correlating highly with human judgment. The results highlight INACIA’s proficiency in handling complex legal tasks, indicating its suitability for augmenting efficiency and judicial fairness within legal systems. The paper also discusses potential enhancements and future applications, positioning INACIA as a model for worldwide AI integration in legal domains.
@@ 1. Introduction
In this paper, we introduce a groundbreaking application of LLMs for administrative case processing and assisted decision-making developed jointly by the Brazilian Federal Court of Accounts (TCU, or Tribunal de Contas da União, in Portuguese) and experts at several Brazilian research organizations. INACIA (Instrução Assistida com Inteligência Artificial, in Portuguese or Instruction Assisted with Artificial Intelligence) is a system that suggests judication directions, which consist of a proposed way for the decision-making authority to address the case. Such directions are fundamentally the responsibility of TCU’s staff auditors. INACIA serves as an assistant in this process, providing analytical support. Its inputs are case filings; outputs are suggestions to the auditor to write the initial instruction [^1] The initial instruction is the document that contains the directions written by the auditor, including the main facts of the case, the legal basis, and the proposals.. INACIA uses a combination of LLMs and search engines to extract information from case documents and to reason through the extracted information.
The main challenge in developing systems like INACIA is evaluating their performance. How can we ensure the quality of the produced text and extracted data? To address this challenge, we construct a validation dataset from TCU’s case filings and propose a methodology that uses LLMs to evaluate the system’s outputs. To our knowledge, this is the first evaluation of this kind in the legal domain. We demonstrate that our methodology is highly correlated with human evaluation and can be used to evaluate the system’s performance. Our results demonstrate that INACIA can produce high-quality text for the initial instruction but also forewarn the need for further research to improve the quality of the system’s outputs.
@@ 2.1. Role and Responsibilities
In most contemporary democracies, a network of institutions exists to regulate, investigate, and potentially sanction individuals occupying elected and non-elected office for improperly using public funds (Speck, 2011). Despite corruption often dominating media attention, mismanagement of public resources makes up the majority of audit findings across Brazil (Ferraz and Finan, 2008). TCU occupies a premier position within the control network of the Brazilian government and is seen as a model institution for auditing practices across the country. [^2] Each state within Brazil, along with a select number of municipalities, maintains its audit court, which typically draws inspiration for its guidelines from the Federal Court of Accounts (TCU).
@@ 2.2. Artificial Intelligence and TCU
Artificial Intelligence (AI) technologies are not new to TCU. Over the past five years, court developers have worked on systems to classify documents by type, summarize case documents, and efficiently search and retrieve case law information to support decisions. Recently, TCU created its ChatGPT-based conversational assistant to handle unstructured, internal information. In 2023, TCU inaugurated an Artificial Intelligence Center (or Núcleo de Inteligência Artificial, in Portuguese), whose goal is to guide and implement responsible, ethical use of AI in the Court.
Acknowledging the rapid evolution of AI and in line with its mission to foster innovation and enhance Public Administration, TCU initiated a series of studies in 2019 on Pre-commercial Procurement (PCP). Subsequently, in 2022, it launched its first “ETEC” (Encomenda Tecnológica, in Portuguese), a type of public procurement that requires Research and Development with technological risk, where the demanded solution extends beyond the current state of the art. Indeed, the envisaged solution is intended to expedite audit case processing by assisting the TCU auditor responsible for the case as effectively as a human specialist in such tasks would be capable of offering. After a competitive and meticulous selection process [^3] Available at https://tcu.gov.br/etec (content in Portuguese), the consortium of companies NeuralMind [^4] https://neuralmind.ai, a pioneer in solutions employing LLMs in Brazil, and TerraNova [^5] https://trnv.com.br/, specialized in jurimetrics, was chosen. INACIA represents the initial proof of concept for the desired outcome of the ETEC initiative. TCU aims to focus auditors’ expertise on the most challenging tasks of case processing while handing over the more routine tasks to the AI system. This strategy is well-aligned with the evidence on how best to incorporate technology for productivity gains and labor output augmentation (Acemoglu and Autor, 2011). The current expectation is that AI will change how TCU handles cases from a sequential, step-by-step approach to a more holistic approach in which filings are evaluated simultaneously and connected to external resources necessary for decision-making from the outset. Effectively, this system would eliminate redundancies and streamline case processing without requiring regulation change.
@@ 2.3. Audit Case Processing
While case proceedings may vary based on the initiation method, case processing is typically a time-intensive task primarily handled by TCU staff. Auditors must review case filings, confirm whether TCU should hear the case, search and append external evidence of wrongdoing (if necessary), check claims against case law and federal statutes, and apply the Court’s jurisprudence. Finally, they prepare a document called Instrução (Portuguese for Instruction) that summarizes the case, the evidence, and the legal analysis and proposes a decision. The instruction is then sent to the court’s judicial panel for final adjudication. Despite the variability in human effort required for each case, it can be reasonably estimated that the review of a single case necessitates approximately 30 hours, incurring a cost of approximately \$1,750 before the case is even presented to the court’s judicial panel. Given an annual average processing rate of about 2,000 cases within the purview of the INACIA project and the textual nature of the task, audit case processing is an ideal candidate for an LLM application that would help expedite review while reducing costs. [^6] For comparison purposes, INACIA processes a case in under 20 minutes and at an initial cost of \$10, subject to further human review. A mere 1 percent reduction in the time required for a human to process a case would suffice to offset the costs associated with INACIA.
@@ 3.1. Overview
Figure 1 summarizes INACIA’s main case processing pipeline, showcasing the holistic nature of case analysis. While there is a direct, sequential flow (i.e., solid arrows) of tasks (i.e., blue boxes) needed to produce the instruction (case suggested decision), the dashed lines show the association (or dependency) between tasks. The lighter grey boxes indicate data objects needed to start processing or to be produced as intermediate or final products for other TCU systems. In the next subsections, we describe each element of this pipeline in detail.
@@ 3.2. Inputs
Whenever Tika fails in text extraction (e.g., it reported zero searchable characters after processing the text), INACIA then deploys the second text extraction solution, which relies on Microsoft’s Azure Form Recognizer service for Optical Character Recognition (OCR) (Microsoft, [n. d.]). Form Recognizer will ingest any unsearchable file and return the extracted text in plain format. We also compared Form Recognizer to other industry-leading OCR frameworks (Tesseract (Smith, 2007), and Google Cloud Vision (Google, [n. d.])) using the high-quality and low-cost criteria. While they are similarly priced, Form Recognizer had higher-quality results and richer extraction metadata (e.g., information on the position and content of tables).777 To assess text extraction from either document type, we evaluated results qualitatively. We manually compared extracted elements from case documents, such as the position of tables, columns, and text, allowing for a more nuanced understanding of the strengths and weaknesses of each tool in practical scenarios. We evaluated a sample of documents from TCU’s database containing all cases and their documents between 2019 and 2022. We pulled out a minimum of five documents per year. Within these documents, we had at least one document of easy, medium, and hard extraction, determined by the number of pages and their content as follows: (1) Easy: up to 10-page documents with minimal tables; (2) Medium: up to 25-page documents with structured data or images; (3) Hard: over 25-page documents with structured and unstructured data elements, images, or handwritten content.
@@ 3.3.1. Basic Information Extraction
Figure 2 illustrates the basic information extraction process. It is a structured and systematic process relying on prompt engineering and retrieval-augmented generation (RAG) techniques (Lewis et al., 2020): we will first feed the LLM a structured list of the information we need to extract from case documents (“Pre-defined form” green box). This list was constructed using older cases and their representação containing the information TCU needs to retain for record-keeping purposes 888 “Basic information” examples are the case ID, case type, plaintiff name, plaintiff ID, etc. . We also send over detailed explanations of what each of the variables is. Finally, we append to our combined prompt from above the first two pages of a case’s representação (“Main document” gray box), which should contain most of the information needed.
@@ 3.3.2. Extraction of Allegations and Requests
An important step in our pipeline is to assess whether a case’s claims and supporting evidence are strong. Often, TCU staff can turn down the review of a case when there is weak evidence of wrongdoing, when the plaintiff cannot straightforwardly articulate their claims that the defendant has broken the law in some shape or form, or even when the plaintiff does not clearly states their request for TCU action. As with other tasks, extracting allegations and actions requested by plaintiffs is a time-consuming task suited for LLM processing.
@@
The data extracted in this step (allegations and requests) are used in three ways: (1) for the admissibility examination, where INACIA reviews cases for elements that would make such case “of public interest” and whether the claims are strong enough to hear the case; (2) for drafting the initial instruction document (cf. Section 3.7); (3) for the Fumus Boni Iuris analysis, which means evaluating a case for the likelihood of success on its merits (cf. Section 3.6).
@@ 3.4. Admissibility Examination
In this task, the agent can compose queries to search for information in case documents and external databases. For example, the agent may want to compose the query quem tem legitimidade para apresentar uma representação ao TCU? (i.e., “who has legitimacy to file a claim to the TCU?”) to search for information about who can file a claim in external databases. If that’s the case, the agent should not be limited to pre-defined queries and keyword-based queries. To facilitate this, we employ NeuralSearchX ((Almeida et al., 2022)), a search engine that integrates the capabilities of the BM25 algorithm and a neural reranker (Nogueira et al., 2020), fine-tuned for Brazilian Portuguese. NeuralSearchX first retrieves documents containing the words from the query and then utilizes its re-ranking mechanism to prioritize the most relevant documents.
@@ 3.5. Periculum in mora Analysis
The analysis, however, is relatively straightforward. We carry out a simple search for exact matches of minuta de contrato (draft contract) in all case documents after a case has been assigned to a TCU staff, except when the case contains fewer than five documents – in which case all documents are considered. This search is essential to flag documents mentioning the draft contract so they are not included in subsequent steps. Two key rules are used to standardize the decision on whether there is Periculum in mora: (1) the existence of an active or signed contract – there is no danger in delay since TCU cannot prevent a contract from coming into force; (2) and (2) the existence of events that prove delay, cancellation, or impediment in contracting – there is no danger in delay because the execution of a contract has been halted. The verification of these rules occurs step-wise. We first verify whether there is any indication that a goods or service provision contract has been signed or is active. The verification is done directly from a keyword search and indirectly by asking the LLM to reason through the documents and verify the existence and status of procurement contracts. The second step is verifying any event delaying, canceling, or suspending a contract from entering into force. Here, we do not perform a direct keyword search for such information; we send documents and prompts straight to the LLM. The outcome of these searches provides an answer regarding the existence of an active contract and any delays or impediments. Finally, the findings are synthesized into a dictionary containing information about the presence of an active contract or delays, along with the document IDs from each search and a suggested decision on whether there is Periculum in mora, which can be either “rejected” or “accepted”.
It is important to note that the role of the LLM is limited to receiving search results and determining the existence of an active contract or impediments and delays. The decision-making flows are hard-coded and not subject to the LLM’s discretion. After the decision is made, the result is passed onto the LLM along with a prompt laying out decision rules, enabling the LLM to draft the text that will be included in the initial instruction document. This process ensures that the conclusions drawn from the Periculum in mora analysis are accurately reflected in the official documentation, preserving the integrity of TCU’s case review process.
@@ 3.7. Recommendations
The INACIA pipeline concludes by generating a paragraph that, in this context, is called recommendations. The recommendations are a set of directions for the decision-making authority to address the case. As the text is the auditor’s responsibility, INACIA only suggests directions for the auditor to consider. This pipeline task is relatively straightforward and recasts the other outputs from each task to construct a cohesive and coherent instruction document to be sent over to the panel of judges who will issue the final ruling on a case. To write case recommendations, INACIA uses the following components:
  • The output from the admissibility examination, i.e., whether the case is admissible and the explanation of such classification.
@@
Each of these inputs is then processed by the LLM, which had been previously instructed with guidelines and examples of recommendations authored by experienced auditors. The LLM generates paragraphs for each recommendations section, ensuring the final text is coherent and accurately reflects the proposals and analyses. The end product is a structured document that is broken down into several sections, each on different aspects of the case, such as the case information, decisions on admissibility, merits, etc.
This modular construction of the final instruction document makes sense both from the legal and the computational sense. From the legal perspective, it allows human reviewers to parse out the document into smaller chunks and use/correct them as they see fit. From the computational side, it means we do not have to reprocess everything in later stages if we need a small correction in one of the earlier stages of INACIA. This approach provides efficiency, especially in cases where, despite a potential flaw in one section, the rest of the information remains valid and useful.